OpenStudy (anonymous):

Which of the following is the solution of log (x - 3)0.001 = -3

OpenStudy (solomonzelman):

\(\large\color{midnightblue}{ \rm log (x - 3)0.001 = -3 }\) \(\large\color{midnightblue}{ \rm log (x - 3) = -3000 }\) \(\large\color{midnightblue}{ \rm log (x - 3) =-3000~log(10) }\) \(\large\color{midnightblue}{ \rm log (x - 3) =log(10^{-3000}) }\) \(\large\color{midnightblue}{ \rm x - 3=10^{-3000} }\) \(\large\color{midnightblue}{ \rm x =3+10^{-3000} }\)

OpenStudy (solomonzelman):

\(\large\color{midnightblue}{ \rm x =3+10^{-3000} }\) \(\large\color{midnightblue}{ \rm x =3+1/(10^{3000} ) }\)

OpenStudy (solomonzelman):

I am not writing all the zeros out. Sorry.

OpenStudy (anonymous):

A) x = 5 B) x = 7 C) x = 13 D) x = 15

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@SolomonZelman i need help

OpenStudy (solomonzelman):

none of them. I showed the solution. Unless you miss-wrote the question.